# A Simple and Effective Pruning Approach for Large Language Models

### v1

@@ 4 Experiments
4.1 Setup
Models. We evaluate Wanda on the LLaMA Touvron et al. (2023) model family, a series of Transformer language models at various parameter levels, often referred to as LLaMA-7B/13B/30B/65B. We apply our pruning method to all four LLaMA models. We note that our approach is not limited to LLaMA, but is applicable to any architectures that consist of many linear layers, including other Transformer-based LLMs. We provide results for other LLM families in Appendix B.
Evaluation. We first measure the performance of pruned networks by their language modeling ability: perplexity computed on a held-out validation set. Following previous works on LLM compression Xiao et al. (2023); Frantar and Alistarh (2023); Frantar et al. (2023), we report the perplexity metric on WikiText Merity et al. (2016) validation set. While perplexity has been shown to be a stable and robust metric for language models Xia et al. (2022a), we also evaluate their zero-shot ability. We use the public evaluation benchmark EleutherAI LM Harness Gao et al. (2021). For zero-shot performance, we evaluate on seven common sense benchmarks: BoolQ Clark et al. (2019), RTE Wang et al. (2018), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2019), ARC Easy and Challenge Clark et al. (2018), and OpenbookQA Mihaylov et al. (2018). We report the accuracy on each benchmark as well as the overall mean accuracy.
Baselines. We compare our method Wanda to two prior pruning methods:
Magnitude pruning (Han et al., 2015) is a simple and strong baseline in which weights are discarded based on their magnitudes. We follow the previous practice of magnitude pruning on Transformers Gale et al. (2019); Sanh et al. (2020), where weights are compared locally among the current linear layer. Magnitude pruning has been demonstrated to outperform many other pruning methods (Gale et al., 2019; Blalock et al., 2020).
SparseGPT (Frantar and Alistarh, 2023) is a second-order pruning method based on solving a layer-wise reconstruction problem. To scale existing second-order based approaches (Frantar et al., 2022) to LLMs, an efficient weight update procedure was proposed that iterates between weight removal and weight update at each layer. To our knowledge, it is the only pruning method so far that has been demonstrated to work with LLMs.
@@ 4.1 Setup
Both Wanda and SparseGPT require calibration data to estimate certain input statistics (see Table 1). To control this variable factor, we use the exact same set of calibration data as SparseGPT, which consists of 128 sequences (2048 tokens each) sampled from the first shard of the C4 training data Raffel et al. (2020).
Sparsity. For all pruning methods, we follow the setup of SparseGPT (Frantar and Alistarh, 2023), where a uniform sparsity is imposed for all layers and there is no subsequent retraining. We skip the first embedding layer and the final classification head, as is common in pruning Transformers Sanh et al. (2020); Frantar and Alistarh (2023). Our primary approach to induce sparsity is through unstructured pruning. Considering the potential need for practical speedup, we also conduct evaluations on structured N:M sparsity Zhou et al. (2021); Hubara et al. (2021). Specifically, we provide comparisons on 4:8 and 2:4 sparsity patterns. The magnitude pruning baseline is extended to structured N:M sparsity in a similar spirit to our method which is described in the previous section.
| LLaMA | ||||||
| Method | Weight Update | Sparsity | 7B | 13B | 30B | 65B |
| Dense | - | 0 | 5.68 | 5.09 | 4.77 | 3.56 |
| Magnitude | ✗ | 50 | 17.29 | 20.21 | 7.54 | 5.90 |
| SparseGPT | ✓ | 50 | 7.22 | 6.21 | 5.31 | 4.57 |
| Wanda | ✗ | 50 | 7.26 | 6.15 | 5.24 | 4.57 |
| Magnitude | ✗ | 4:8 | 16.84 | 13.84 | 7.62 | 6.36 |
| SparseGPT | ✓ | 4:8 | 8.61 | 7.40 | 6.17 | 5.38 |
| Wanda | ✗ | 4:8 | 8.57 | 7.40 | 5.97 | 5.30 |
| Magnitude | ✗ | 2:4 | 42.13 | 18.37 | 9.10 | 7.11 |
| SparseGPT | ✓ | 2:4 | 11.00 | 9.11 | 7.16 | 6.28 |
| Wanda | ✗ | 2:4 | 11.53 | 9.58 | 6.90 | 6.25 |
@@ 4.2 Language Modeling
Unstructured Sparsity. For each of the LLaMA models, we prune it to unstructured 50% sparsity for all methods. Results are shown in Table 2. Without any weight update, Wanda outperforms the established pruning approach of magnitude pruning by a large margin. For instance, for LLaMA-7B, Wanda is able to find sparse networks with a perplexity of 7.26, significantly better than the magnitude pruning baseline 17.29. Our method also performs on par with or in most cases better than the prior reconstruction-based method SparseGPT.
Structured N:M Sparsity. We now turn our eyes to structured N:M sparsity. Note that N:M sparsity by definition is a more restrictive sparsity pattern than unstructured sparsity. Thus, it is expected that such structured sparsity patterns would lead to worse results. Results for structured 4:8 and 2:4 sparsity are shown in the lower parts of Table 2. We can see that Wanda can be easily generalized to structured N:M sparsity. Across 4:8 and 2:4 sparsity, our method consistently finds highly effective sparse sub-networks, outperforming baseline approaches in most cases, especially for larger models (e.g. 30B and 65B).
Remark. Note that in certain cases, Wanda obtains results slightly better than but somewhat close to those of SparseGPT. This may be related to the fact that our pruning metric shares an implicit connection to the OBS reconstruction error in Equation 3, as identified and explained in Section 3. While perplexity improvements in these cases over SparseGPT may not be considered substantial, Wanda is much simpler and more computationally efficient.
Varying Sparsity Levels. We conduct experiments with varying levels of sparsity for unstructured pruning, the results of which are depicted in Figure 2. It can be seen that Wanda and SparseGPT show similar trends of perplexity increase as the sparsity level gets higher. However, magnitude pruning displays a considerably more severe degradation trend.
4.3 Zero-Shot Tasks
We evaluate pruned models on downstream zero-shot tasks via prompting. Results are summarized in Table 3, where models are pruned to unstructured 50% sparsity. Averaging the accuracy over the 7 tasks under consideration, the Wanda method is able to identify effective pruned networks, showing competitiveness with baseline methods. Note that as the model gets larger in size, the accuracy drop compared to the original dense model keeps getting smaller. For task-wise performance, we observe that there are certain tasks where our approach Wanda gives consistently better results across all LLaMA models, i.e. HellaSwag, ARC-c and OpenbookQA, while among the remaining tasks, there is not a fixed superior between the two methods.
| 7B | Dense | 71.7 | 53.4 | 58.3 | 68.0 | 67.7 | 38.6 | 28.0 | 55.1 |
| Magnitude | 53.5 | 56.3 | 44.6 | 54.5 | 53.7 | 32.3 | 21.6 | 45.2 | |
| SparseGPT | 71.5 | 56.8 | 52.7 | 65.8 | 64.3 | 36.0 | 25.0 | 53.2 | |
| Wanda | 70.3 | 53.3 | 52.9 | 64.7 | 64.1 | 37.0 | 26.4 | 52.7 | |
| 13B | Dense | 68.3 | 65.3 | 60.8 | 70.0 | 73.6 | 44.0 | 30.6 | 58.9 |
| Magnitude | 62.1 | 45.8 | 44.3 | 58.9 | 49.8 | 33.1 | 27.4 | 45.9 | |
| SparseGPT | 66.7 | 52.0 | 55.1 | 70.9 | 66.9 | 39.2 | 26.0 | 53.8 | |
| Wanda | 67.1 | 61.0 | 56.6 | 71.7 | 68.4 | 41.7 | 28.0 | 56.4 | |
| 30B | Dense | 66.9 | 61.4 | 64.8 | 72.4 | 75.3 | 46.9 | 29.4 | 59.6 |
| Magnitude | 66.6 | 53.4 | 49.7 | 63.4 | 68.8 | 40.0 | 26.2 | 52.6 | |
| SparseGPT | 71.0 | 61.4 | 61.1 | 72.0 | 73.9 | 46.0 | 31.2 | 59.5 | |
| Wanda | 70.3 | 66.8 | 62.3 | 71.1 | 74.8 | 46.5 | 32.4 | 60.6 | |
| 65B | Dense | 81.8 | 71.8 | 65.2 | 76.9 | 75.4 | 47.2 | 36.4 | 65.0 |
| Magnitude | 79.8 | 66.8 | 62.5 | 70.3 | 70.4 | 46.7 | 31.4 | 61.1 | |
| SparseGPT | 81.2 | 70.4 | 62.5 | 74.1 | 74.8 | 44.7 | 32.2 | 62.8 | |
| Wanda | 82.2 | 69.1 | 64.5 | 73.9 | 74.6 | 45.9 | 32.8 | 63.3 |
4.4 Pruning Speed
@@
The theoretical computational complexity of Wanda is lower than SparseGPT (Table 1). Here we compare their empirical pruning speed. We measure the total pruning time only (excluding the forward pass process shared by both methods) on NVIDIA A6000 GPUs. We present the results in Table 4. Wanda incurs negligible time overhead compared with SparseGPT. The fast speed is particularly useful when pruning needs to be performed on a real-time basis, e.g., processing different samples with different sparsity levels.
@@
4.5 Ablation Study
We study three aspects of Wanda, aiming to answer the following questions: How does the choice of pruning configurations, i.e. pruning metric and pruning granularity, impact the performance of pruned models? How robust is our pruning method to calibration samples? Can Wanda be integrated with the OBS based weight update as used in SparseGPT? For the following ablation studies, we prune the LLaMA-7B model to unstructured 50% sparsity and report the perplexity on WikiText.
Pruning Configuration. Wanda differs from previous methods in both the pruning metric and the pruning granularity. We conduct experiments on a variety of pruning metrics and grnularities to ablate their effectiveness. The three pruning metrics can be found in Table 1. SparseGPT adopts a local granularity level inside a layer, where weights connected to 128 consecutive input channels form a group. Wanda groups weights connected with a single output channel. Therefore, we ablate two blocksize options (128 and 1) and the input/output choice. For simplicity, we use (input/output, blocksize) to denote the local pruning granularity level, e.g., (input, 1). For this experiment, we do not perform weight update for SparseGPT to focus on metric and granularity.
| Pruning Granularity | |||||
@@ 4.5 Ablation Study
The results are shown in Table 5. Wanda’s default configuration delivers the best pruned model (perplexity 7.26). Interestingly, for the magnitude metric, comparing weights of the same input neuron (input, 1) yields a perplexity of 8.86, significantly better than other granularity options. Three methods also produce equivalent pruning results as under this granularity – the input is the same, thus weight ranking only depends on weight magnitude. This finding further highlights the importance of using a proper pruning granularity for pruning LLMs, even for the classical magnitude pruning.
Robustness to Calibration Samples. Our method relies on a set of calibration data to estimate input statistics of each layer. It is thus important to understand how the construction of calibration data affects the final pruned networks. We sample 5 sets of calibration data with the same size (128) from different shards of C4 training data. We find that pruned networks obtained from 5 calibration sets have a perplexity of 7.21, 7.29, 7.23, 7.25, 7.25, respectively. The average perplexity is 7.25, close to the 7.26 we reported previously. This indicates the performance does not depend too heavily on the exact selection of calibration data.
@@
We then vary the number of calibration samples by selecting different sample sizes ranging between 1 and 256. Results are summarized in Figure 3. We see a clear difference in trend as the size of calibration data changes, where Wanda is much more robust when the calibration samples are few. Notably, even with a single sample, pruned networks obtained by our approach have a perplexity of 7.66. This may be because input norm statistics could be much easier to estimate than the full inverse hessian of the local layer-wise reconstruction problem.
Weight Update. We test if we could obtain better pruned models by incorporating the OBS weight update process proposed in SparseGPT Frantar and Alistarh (2023). We apply the OBS weight update procedure used in SparseGPT on the kept weights produced by both magnitude pruning and our approach Wanda. We evaluate the perplexity on both the calibration data and the validation data from WikiText.
Results are summarized in Table 6. The result on magnitude pruning does validate the effectiveness of the weight update procedure, where it improves perplexity significantly on both data splits. However, when the pruned model is obtained by our method Wanda, it no longer brings improvement but leads to an increase of perplexity (from 7.26 to 7.32) on validation data, suggesting that in this case the updated model may overfit to the calibration samples.
5 Analysis
In this section, we first study whether Wanda can be a general pruning approach beyond LLMs. Next, we show exploratory results on using parameter efficient fine-tuning techniques to recover performance drop during the pruning procedure.
5.1 Generality beyond Pruning LLMs
While the main focus of this work is on pruning LLMs, the surprising effectiveness of Wanda leads to a clear question: how would Wanda perform against magnitude pruning on tasks where the latter has been widely used? We thus conduct a study on ImageNet-1K (Deng et al., 2009), a standard image classification task where magnitude pruning has been extensively studied and accepted as a strong baseline Gale et al. (2019); Blalock et al. (2020).
We consider two modern vision architectures: ConvNeXt Liu et al. (2022) and Vision Transformer (ViT) Dosovitskiy et al. (2021). We choose these two architectures mainly for two reasons: first, as LLMs are based on Transformers Vaswani et al. (2017), we would like to test if our observations on LLMs so far still hold on Transformers for other tasks; second, as we are evaluating on image classification, we are interested in examining how they work with ConvNet models, with ConvNeXt being a representative architecture.
We use two ImageNet-1K pretrained models: ConvNeXt-B and ViT-B, which have a top-1 accuracy of 83.8% and 81.8% Touvron et al. (2021) respectively. We prune the linear layers only (for ConvNeXt, this includes equivalent 11 convolution layers). We find that for both models, pruning uniformly on a per layer basis is consistently better than a per output basis (detailed results are shown in Appendix B). We thus compare our proposed metric and the standard magnitude metric by pruning on a per layer basis. Results are shown in Figure 4. Our novel pruning metric leads to better results than magnitude pruning, especially at high sparsities (e.g. 70%, 80%). However, we find differences in the accuracy of pruned networks can easily be mitigated via retraining for a few epochs, indicating that Wanda might be more suitable for cases where retraining is not computationally feasible, such as in LLMs.
5.2 LoRA Fine-tuning
We explore using parameter efficient fine-tuning (PEFT) techniques to recover performance of pruned LLM models. We use a popular PEFT method LoRA Hu et al. (2021), which has been widely adopted for task specific fine-tuning of LLMs. However, here we are interested in recovering the performance loss of LLMs during pruning, thus we perform a more general “fine-tuning” where the pruned networks are trained with an autoregressive objective on C4 dataset. We enforce a limited computational budget (1 GPU and 5 hours). We find that we are able to restore performance of pruned LLaMA-7B (unstructured 50% sparsity) with a non-trivial amount, reducing zero-shot WikiText perplexity from 7.26 to 6.87. The additional parameters introduced by LoRA is only 0.06%, leaving the total sparsity level still at around 50% level.

### v3

@@ 4 Experiments
Models and Evaluation. We evaluate Wanda on the two most widely adopted LLM model families: LLaMA 7B/13B/30B/65B (Touvron et al., 2023a) and LLaMA-2 7B/13B/70B (Touvron et al., 2023b) (LLaMA-2 34B is not released). Results for prior LLM families can be found in Appendix B. We measure the performance of pruned models on zero-shot tasks and language modeling. For zero-shot evaluation, we use seven tasks from EleutherAI LM Harness (Gao et al., 2021). Following previous works on LLM compression (Xiao et al., 2023; Frantar & Alistarh, 2023), we evaluate the perplexity on the held-out WikiText (Merity et al., 2016) validation set.
Baselines. We compare Wanda with two prior pruning approaches. Magnitude pruning (Han et al., 2015) is a simple and strong baseline in which weights are discarded based on their magnitudes. SparseGPT (Frantar & Alistarh, 2023) is a second-order pruning method for LLMs, based on solving a layer-wise reconstruction problem. In Appendix C, we compare with additional pruning methods.
@@
Both Wanda and SparseGPT require calibration data to estimate input statistics (see Table 1). To control this variable factor, we use the exact same set of calibration data as SparseGPT, which consists of 128 sequences with context length size sampled from C4 training set (Raffel et al., 2020). In Appendix D.1, we provide additional analysis on the number of calibration samples.
Sparsity. For all pruning methods, we focus on pruning the linear layers (skipping the first embedding layer and the final classification head), which account for around 99 of the total LLM parameters. We impose a uniform sparsity for all linear layers. We evaluate three types of sparsity: unstructured sparsity, structured 4:8 and 2:4 sparsities. The magnitude pruning baseline is extended to structured N:M sparsity in a similar spirit to our method, as described in the previous section.
| LLaMA | LLaMA-2 | ||||||||
| Method | Weight Update | Sparsity | 7B | 13B | 30B | 65B | 7B | 13B | 70B |
| Dense | - | 0 | 59.99 | 62.59 | 65.38 | 66.97 | 59.71 | 63.03 | 67.08 |
| Magnitude | ✗ | 50 | 46.94 | 47.61 | 53.83 | 62.74 | 51.14 | 52.85 | 60.93 |
| SparseGPT | ✓ | 50 | 54.94 | 58.61 | 63.09 | 66.30 | 56.24 | 60.72 | 67.28 |
| Wanda | ✗ | 50 | 54.21 | 59.33 | 63.60 | 66.67 | 56.24 | 60.83 | 67.03 |
| Magnitude | ✗ | 4:8 | 46.03 | 50.53 | 53.53 | 62.17 | 50.64 | 52.81 | 60.28 |
| SparseGPT | ✓ | 4:8 | 52.80 | 55.99 | 60.79 | 64.87 | 53.80 | 59.15 | 65.84 |
| Wanda | ✗ | 4:8 | 52.76 | 56.09 | 61.00 | 64.97 | 52.49 | 58.75 | 66.06 |
| Magnitude | ✗ | 2:4 | 44.73 | 48.00 | 53.16 | 61.28 | 45.58 | 49.89 | 59.95 |
| SparseGPT | ✓ | 2:4 | 50.60 | 53.22 | 58.91 | 62.57 | 50.94 | 54.86 | 63.89 |
| Wanda | ✗ | 2:4 | 48.53 | 52.30 | 59.21 | 62.84 | 48.75 | 55.03 | 64.14 |
4.1 Zero-Shot Tasks
Comparison with Baselines. In Table 2, we show the mean zero-shot accuracies on 7 zero-shot tasks of pruned LLaMA and LLaMA-2 models. We refer the reader to Appendix D for task-wise performance. Across both unstructured and structured sparsities, Wanda outperforms the well-established magnitude pruning approach by a large margin, while also rivals with the previous best approach SparseGPT. Given that no fine-tuning takes place, there is a noticeable gap between sparse pruned LLMs and the original dense LLMs. However, as the model size increases, this accuracy gap diminishes. Remarkably, unstructured 50 sparse LLaMA-65B and LLaMA-2-70B is able to match the zero-shot accuracies of their dense counterparts.
Large Sparse vs. Small Dense. It might be of interest to some readers on the comparison between large sparse LLMs and small dense LLMs with similar parameter counts. For zero-shot performance, we find the trend differs across the types of sparsity. For unstructured sparsity, large sparse LLMs are often better than small dense LLMs on zero-shot performance: unstructured 50 sparse LLaMA-65B (66.67) outperforms dense LLaMA-30B (65.38); unstructured 50 sparse LLaMA-2-13B (60.83) outperforms dense LLaMA-7B (59.71). Intriguingly, this gap is much larger for few-shot tasks (see Appendix D). For structured sparsity, the trend is reversed: without any fine-tuning, large sparse LLMs have worse zero-shot performance than small dense LLMs in general.
@@ 4.2 Language Modeling
In Table 3, we report the perplexity of pruned LLaMA and LLaMA-2 models. For robustness analysis under random sampling of the calibration data, see Appendix D.
@@
Without any weight update, Wanda outperforms the established pruning approach of magnitude pruning by a large margin. For instance, for LLaMA-7B, Wanda is able to find sparse networks with a perplexity of 7.26, significantly better than the magnitude pruning baseline 17.29. This result suggests that exact and effective sparse sub-networks exist for LLMs. For unstructured 50% sparsity, Wanda performs on par with the prior best approach SparseGPT. We provide results for higher sparsity levels (60% and 80%) in Appendix D. The comparison between Wanda and SparseGPT is mixed for structured sparsity. On smaller models (e.g., 7B), SparseGPT outperforms Wanda on 2:4 sparsity. Wanda is more favorable for larger models, e.g., LLaMA-30B (2:4 and 4:8) and LLaMA-2-70B (2:4).
| LLaMA | LLaMA-2 | ||||||||
| Method | Weight Update | Sparsity | 7B | 13B | 30B | 65B | 7B | 13B | 70B |
| Dense | - | 0 | 5.68 | 5.09 | 4.77 | 3.56 | 5.12 | 4.57 | 3.12 |
| Magnitude | ✗ | 50 | 17.29 | 20.21 | 7.54 | 5.90 | 14.89 | 6.37 | 4.98 |
| SparseGPT | ✓ | 50 | 7.22 | 6.21 | 5.31 | 4.57 | 6.51 | 5.63 | 3.98 |
| Wanda | ✗ | 50 | 7.26 | 6.15 | 5.24 | 4.57 | 6.42 | 5.56 | 3.98 |
| Magnitude | ✗ | 4:8 | 16.84 | 13.84 | 7.62 | 6.36 | 16.48 | 6.76 | 5.54 |
| SparseGPT | ✓ | 4:8 | 8.61 | 7.40 | 6.17 | 5.38 | 8.12 | 6.60 | 4.59 |
| Wanda | ✗ | 4:8 | 8.57 | 7.40 | 5.97 | 5.30 | 7.97 | 6.55 | 4.47 |
| Magnitude | ✗ | 2:4 | 42.13 | 18.37 | 9.10 | 7.11 | 54.59 | 8.33 | 6.33 |
| SparseGPT | ✓ | 2:4 | 11.00 | 9.11 | 7.16 | 6.28 | 10.17 | 8.32 | 5.40 |
| Wanda | ✗ | 2:4 | 11.53 | 9.58 | 6.90 | 6.25 | 11.02 | 8.27 | 5.16 |
4.3 Speedup
@@
Pruning Speed. The theoretical computational complexity of Wanda is lower than SparseGPT (Table 1). Here we compare their empirical pruning speed. Specifically, we measure the accumulated time for computing the pruning metric at each layer (excluding the forward pass process shared by both methods) on NVIDIA A6000 GPUs. Results are shown in Table 4. Wanda incurs negligible time overhead relative to SparseGPT. The fast speed of Wanda is particularly useful when pruning needs to be performed on a real-time basis, e.g., training sparse models from scratch (Evci et al., 2020) and finding the optimal sparsity (Jin et al., 2022).
Inference Speed. We evaluate the inference speedup for structured 2:4 sparsity on NVIDIA A6000 GPUs. Following the evaluation setup of Frantar & Alistarh (2023), we measure the latency of matrix multiplication in linear layers. We perform simulation analysis using the high-performance GEMM kernel in NVIDIA CUTLASS library. Results for LLaMA-65B (batch size of 1) can be found in Table 5. Structured 2:4 sparsity is able to bring notable inference speedup (around 1.6) for linear layers in LLMs. For end to end latency, we observe a speedup of 1.24 on LLaMA-7B (251ms as compared to 312ms). Last, we emphasize that the inference speedup is not unique to our pruning method but is delivered by the inherent power of sparsity for speeding up computation.
@@
| LLaMA Layer | Dense | 2:4 | Speedup |
| q/k/v/o_proj | 3.49 | 2.14 | 1.63 |
| up/gate_proj | 9.82 | 6.10 | 1.61 |
| down_proj | 9.92 | 6.45 | 1.54 |
5 Analysis
We study several aspects of Wanda to better understand its effectiveness in pruning LLMs. We use the LLaMA-7B model and prune to unstructured 50 sparsity, unless otherwise specified.
Fine-tuning. We study how fine-tuning could recover the performance drop of pruned LLMs, as observed in the previous section. We investigate two strategies for fine-tuning LLMs: LoRA (Hu et al., 2021) fine-tuning and full parameter dense fine-tuning. Fine-tuning is conducted on C4 training dataset and the objective is the pre-training auto-regressive loss. The pruned mask is kept fixed during fine-tuning. We fine-tune pruned LLaMA-7B with all three types of sparsities: unstructured 50, structured 4:8 and 2:4. Table 6 summarizes the results for mean zero-shot accuracies and perplexity after fine-tuning Wanda pruned LLaMA-7B models. See Appendix D for task-wise performance.
| Evaluation | Dense | Fine-tuning | 50% | 4:8 | 2:4 |
| Zero-Shot | 59.99 | ✗ | 54.21 | 52.76 | 48.53 |
| LoRA | 56.53 | 54.87 | 54.46 | ||
| Full | 58.15 | 56.65 | 56.19 | ||
| Perplexity | 5.68 | ✗ | 7.26 | 8.57 | 11.53 |
| LoRA | 6.84 | 7.29 | 8.24 | ||
| Full | 5.98 | 6.63 | 7.02 |
LoRA Fine-tuning. We enforce a limited computational budget (1 GPU and 12 hours). The low rank () adapter is applied on the query and value projection matrices in attention layers. For LLaMA-7B, LoRA introduces only around 0.06% additional parameters, leaving the total sparsity level still around 50%. With LoRA fine-tuning, we are able to restore the performance of pruned LLMs by a non-trivial amount. One notable instance is that LoRA fine-tuning improves the zero-shot performance of structured 2:4 sparse LLaMA-7B from 48.53 to 54.46, outperforming the original unstrucutred 50 sparse LLaMA-7B (54.21).
Full Parameter Fine-tuning. We conduct full parameter dense fine-tuning. We enforce a limited computational budget (4 GPU and 3 days). Compared to LoRA fine-tuning, full parameter dense fine-tuning is able to mitigate the gap between pruned LLMs and dense LLMs even further. For unstructured 50 sparsity, full parameter fine-tuning could improve pruned LLaMA-7B from 54.21 to 58.15 in terms of zero-shot accuracy, close to that of dense LLaMA-7B (59.99).
Pruning Configuration. Wanda differs from previous methods in both the pruning metric and the comparison group. We conduct ablation experiments to better understand their impact. The three pruning metrics can be found in Table 1. SparseGPT adopts a local comparison group inside a layer, where weights connected to 128 consecutive input channels form a group. Wanda groups weights connected with a single output channel. Therefore, we ablate two blocksize options (128 and 1) and the input/output choice. For simplicity, we use (input/output, blocksize) to denote each local comparison group, e.g., (input, 1). For this experiment, we do not perform the weight update procedure in SparseGPT to focus on the pruning configuration.
| Comparison Group | |||||
@@ 5 Analysis
The results are shown in Table 7. We refer the reader to Appendix A for analysis on image classifiers and Appendix D for analysis on previous LLMs. The default pruning configuration of Wanda delivers the best pruned model (perplexity 7.26). Interestingly, for the magnitude metric, comparing weights of the same input neuron (input, 1) yields a perplexity of 8.86, significantly better than other grouping options. Three methods also produce equivalent pruning results as under this comparison group – the input is the same, thus weight ranking only depends on weight magnitude. This finding further highlights the importance of using a proper comparison group for pruning LLMs, even for the classical magnitude pruning approach.
@@
Robustness to Calibration Samples. We vary the number of calibration samples by selecting different sample sizes ranging between 1 and 256. Results are summarized in Figure 2. We see a clear difference in trend as the size of calibration data changes, where Wanda is much more robust when there are few calibration samples. Notably, even with a single sample, pruned networks obtained by Wanda have a perplexity of 7.66. This may be because input norm statistics could be much easier to estimate than the full inverse hessian of the local layer-wise reconstruction problem.
Weight Update. We characterize the conditions under which the weight update process in SparseGPT can improve the effectiveness of pruning LLMs. We experiment with two ways of applying weight update: sequential and iterative. A sequential update means that at each layer, the full pruned mask is first computed and weight update is performed on the remaining weights. An iterative update means that the pruning and weight update steps proceed iteratively within each layer. SparseGPT adopts an iterative update procedure every 128 input channels, as it was found to give more accurate results.
| Pruning Configuration | Weight Update | Sparsity | |||
| Pruning Metric | Comparison Group | 50% | 4:8 | 2:4 | |
| Magnitude: | layer | ✗ | 17.59 | 16.84 | 42.13 |
| layer | Sequential | 12.56 | 13.37 | 21.36 | |
| (input, 128) | Iterative | 26.77 | 36.98 | 47.61 | |
| Wanda : | (output, 1) | ✗ | 7.26 | 8.57 | 11.53 |
| (output, 1) | Sequential | 7.32 | 8.59 | 10.89 | |
| (input, 128) | Iterative | 7.26 | 8.68 | 11.43 | |
Effects of the weight update on magnitude pruning and Wanda are summarized in Table 8. We study these two pruning methods because they do not involve any weight update by default. An iterative update changes the comparison group for unstructured pruning, which we denote in the table as (input, 128). We make several interesting observations:
For all considered sparsities, weight update can improve magnitude pruning by a large margin.
For unstructured 50 and 4:8 sparsities, weight update does not bring any improvement to Wanda.
For 2:4 sparsity, the improvement (from 11.53 to 10.89) is marginal. Note that the best 2:4 sparse model (10.89) we obtained here is better than that obtained by SparseGPT (11.00 in Table 3).
Last, we examine an extreme sparsity level (70), where weight update can improve Wanda from 84.50 to 29.65. However, the best pruned model (29.65) lags far behind the dense LLaMA-7B (5.68).